# Efficient simulation of the SABR model

### v1

@@ 4.1 CEV approximation of the conditional price
$\bar{F}_{t}^{h}\approx F_{t}\exp\left(\frac{\rho(\sigma_{t+h}-\sigma_{t})}{\nu F_{t}^{\beta_{\ast}}}-\frac{\rho^{2}\sigma_{t}^{2}hI_{t}^{h}}{2F_{t}^{2\beta_{\ast}}}\right).$ (11)
Proof.
We provide the rationales for each component of the CEV approximation. First, with regard to distribution, the choice of $\mathcal{CEV}_{\beta}$ is a natural extension of the $\nu=0$ and $\rho=0$ cases, where the forward price is distributed by $\mathcal{CEV}_{\beta}$ unconditionally and conditionally, respectively. Given that we choose the CEV distribution as the approximation, we show how to select the mean and variance. The variance, $\rho_{\ast}^{2}\sigma_{t}^{2}hI_{t}^{h}$, is also a natural choice, since it is consistent with all special cases in Section 2.3. The variance in the $\rho=0$ case is merely the special cases of $\rho_{\ast}^{2}\sigma_{t}^{2}hI_{t}^{h}$ when $\rho_{\ast}=1$.
Lastly, we choose the conditional expectation $\bar{F}_{t}^{h}$ as the mean of the CEV distribution. Note that it is analytically intractable to obtain the exact conditional expectation under the original SABR dynamics when $\rho\neq 0$. Instead, we approximate the SABR dynamics in the interval $[\,t,t+h]$ by the lognormal SABR model:
$\frac{\mathrm{d}F_{s}}{F_{s}}=\frac{\sigma_{s}}{F_{s}^{\beta_{\ast}}}\,\mathrm{d}W_{s}\approx\frac{\sigma_{s}}{F_{t}^{\beta_{\ast}}}\,\mathrm{d}W_{s}\quad\text{for}\quad t\leq s\leq t+h,$ (12)
where the stochastic volatility $\sigma_{s}$ is replaced by $\sigma_{s}/F_{t}^{\beta_{\ast}}$. Under this approximation, the conditional expectation in Eq. (11) follows from $\beta=1$ case in Eq. (7). Note that the lognormal SABR approximation is too loose to be used for sampling $F_{t+h}$. We only use it for approximating $\bar{F}_{t}^{h}$. Most importantly, our choice of $\bar{F}_{t}^{h}$ ensures the martingale property of $F_{t}$. Based on Remark 1, we observe
@@
when $F_{t+h}$ is sampled according to this proposition. The martingale condition significantly improves the accuracy of our simulation algorithm (see numerical results illustrated in Figure 3). ∎
4.2 Sampling conditional price from a CEV distribution
The availability of analytic distribution function does not guarantee an efficient sampling algorithm of the random variable. The slow root-finding step is often used to invert the distribution function, and it has been the case with some existing SABR simulation methods (e.g., Cai et al. (2017)). Our SABR simulation algorithm would be completed when our CEV approximation in Proposition 4 is paired with an efficient sampling algorithm of the CEV distribution. For the purpose, we adopt the algorithms of Makarov and Glew (2010) and Kang (2014) for sampling the CEV distribution exactly. Despite its simplicity and efficiency, the algorithm is not widely known in the literature.
Before further discussion, let us present probability distributions that are related to the CEV distribution, and review several related analytic properties.
@@ 4.2 Sampling conditional price from a CEV distribution
Let ${\chi^{2}}(\delta,r)$ be the NCX2 random variable with degree of freedom $\delta$ and noncentrality $r$. The PDF and CDF of ${\chi^{2}}(\delta,r)$ are respectively expressed by
@@
The central chi-squared (i.e., $r=0$) distribution is a gamma distribution. Let $\mathcal{G}(\alpha)$ be a gamma random variable with shape parameter $\alpha$ and unit scale parameter, the two random variables are related by
@@
The PDF and CDF of $\mathcal{G}(\alpha)$ are respectively expressed by  [^3] The CDF, $P_{\mathcal{G}}(x;\alpha)$, is the scaled lower incomplete gamma function $\gamma(\alpha,x)$: $P_{\mathcal{G}}(x;\alpha)=\frac{\gamma(\alpha,x)}{\Gamma(\alpha)}\quad\text{where}\quad\gamma(\alpha,x)=\int_{0}^{x}t^{\alpha-1}e^{-t}\,\mathrm{d}t.$
@@
Definition 4 (Shifted Poisson random variable).
The shifted Poisson (SP) variate with intensity $\lambda$ and shift $\alpha$, $N\sim\mathcal{SP}(\lambda,\alpha)$, takes non-negative integer values with mass probability function:
@@
Next, we discuss the connection of the above defined distributions to the CEV distribution. It is well known that the transition density of $F_{T}$ is closely related to the NCX2 distribution (Schroder, 1989; Cox, 1996).
Proposition 5.
Define $z_{t}$ be a transformation of the CEV process, $F_{t}$, in Eq. (8) by the function $z(\cdot)$:
$z_{t}:=z(F_{t})\quad(0\leq t\leq T)\quad\text{and}\quad z(y):=\frac{y^{2\beta_{\ast}}}{\beta_{\ast}^{2}\sigma_{0}^{2}T}.$ (17)
@@
Since the location variable $z$ plays the role of the noncentrality parameter, the distribution of $z_{T}$ is not a NCX2 distribution. The complementary CDF of $F_{T}>0$ and the probability of absorption (i.e., mass at zero) are respectively given by
$\mathrm{Prob}(F_{T}>y)$ $=\mathrm{Prob}(z_{T}>z(y))=P_{\chi^{2}}\left(z_{0};\frac{1}{\beta_{\ast}},z(y)\right),$ (18)
$\mathrm{Prob}(F_{T}=0)$ $=\mathrm{Prob}(z_{T}=0)=1-P_{\chi^{2}}\left(z_{0};\,\frac{1}{\beta_{\ast}},0\right)=1-P_{\mathcal{G}}\left(\frac{z_{0}}{2};\,\frac{1}{2\beta_{\ast}}\right).$ (19)
@@
Proposition 6 (Mixture gamma representation of the CEV distribution (Makarov and Glew, 2010, § 3.3)).
Under the CEV model with absorbing boundary condition, given $z_{T}>0$ ($F_{t}$ is not absorbed at the origin until time $T$), the transition from $z_{0}$ to $z_{T}$ can be sampled by an SP-mixture gamma distribution:
$z_{T}\sim 2\,\mathcal{G}(N+1)\quad\text{where}\quad N\sim\mathcal{SP}\left(\frac{z_{0}}{2},\frac{1}{2\beta_{\ast}}\right)$
@@
$=P_{\mathcal{G}}\left(\frac{z_{0}}{2};\,\alpha\right)\sum_{k=0}^{\infty}P_{\mathcal{SP}}\left(k;\frac{z_{0}}{2},\alpha\right)\cdot\frac{1}{2}f_{\mathcal{G}}\left(\frac{z_{T}}{2};k+1\right).$
@@
Proposition 7 (Sampling an SP variable (Kang, 2014, Algorithm 3)).
The SP variable, $N\sim\mathcal{SP}(\lambda,\alpha)$, can be sampled by a gamma-mixture Poisson variable:
$N\sim\mathcal{POIS}(\lambda-X)\quad\text{where}\quad X\sim\mathcal{G}(\alpha)\quad\text{conditional on}\quad X<\lambda.$
Proof.
For proof, see Kang (2014). ∎
Combining Propositions 6 and 7 leads to an exact simulation algorithm for sampling the CEV distribution. Moreover, Proposition 7 can be further optimized in the context of the CEV simulation. This is the reason we select ‘Algorithm 3’ among the three algorithms in Kang (2014). As this advantage is not mentioned in Kang (2014), we state explicitly it as a remark:
@@
If simulating $N\sim\mathcal{SP}(\lambda,\alpha)$ is the final goal, $X<\lambda$ can be obtained by repeatedly sampling $X$ until $X<\lambda$. If $N$ is an intermediate variable for the CEV simulation in Proposition 6, however, the repeated sampling of $X$ is unnecessary since
@@
Algorithm 1 Exact simulation of the CEV distribution $F_{T}$ in Eq. (8)
The CEV sampling algorithm can be applied to the conditional distribution of $F_{t+h}$ in Proposition 4. We need to sample $z_{t+h}=z(F_{t+h})$ from $z_{t}=z(\bar{F}_{t}^{h})$, where the transformation $z(\cdot)$ is modified in the context of Proposition 4:
@@
       $I_{t}^{h}\leftarrow$ Proposition 3 ;
       $\bar{F}_{t}^{h}\leftarrow$ Eq. (11) ;
@@
4.3 Comparison to Islah (2009)’s approximation used in other simulation schemes
While earlier SABR simulation methods show various approach for sampling $I_{t}^{h}$ (Step 1), almost all of them (Chen et al., 2012; Cai et al., 2017; Leitao et al., 2017a, b; Grzelak et al., 2019; Cui et al., 2021; Kyriakou et al., 2023) are based on Islah (2009)’s approximation for sampling $F_{t+h}$ (Step 2). We compare our approach to Islah (2009)’s in this section.
Proposition 8 (Islah (2009)’s approximation).
For $10<\beta<$ 1, conditional on $\sigma_{t+h}$ and $I_{t}^{h}$ (and the filtration up to time $t$), the distribution of $F_{t+h}$ is approximated by
@@ 4.3 Comparison to Islah (2009)’s approximation used in other simulation schemes
$z^{\prime}_{t}=\frac{1}{\beta_{\ast}^{2}\rho_{\ast}^{2}\sigma_{t}^{2}hI_{t}^{h}}\left[F_{t}+\frac{\beta_{\ast}\rho}{\nu}(\sigma_{t+h}-\sigma_{t})\right]^{2}\quad\text{and}\quad z(y)=\frac{y^{2\beta_{\ast}}}{\beta_{\ast}^{2}\rho_{\ast}^{2}\sigma_{t}^{2}hI_{t}^{h}}.$
Proof.
For the ideas behind this approximation, see Islah (2009, § 2.2), Chen et al. (2012, Result 2.4), or Cui et al. (2021, Appendix B.1)
Islah’s approximation is compared with the probability under our CEV approximation in Proposition 4:
@@
The key difference between the two is that Islah’s approximation uses the modified degree of freedom, $(1-\beta_{\ast}\rho^{2})/(\beta_{\ast}\rho_{\ast}^{2})$, in Eq. (20), and it is not equal to $1/\beta_{\ast}$ (except $\rho=0$). However, the transformation $z(\cdot)$ uses $\beta_{\ast}$ inconsistently. On the other hands, our CEV approximation uses $1/\beta_{\ast}$, which is the same degree of freedom for the CEV distribution. This means that $F_{t+h}$ under Islah’s approximation is not a CEV distribution, which leads to two drawbacks discussed in the next two remarks.
@@
In order to preserve the martingale property, Leitao et al. (2017a, b) artificially add an additive adjustment, commonly called the martingale correction, to the simulated samples of $F_{t+h}$. This ad-hoc correction is inferior to our approach where the martingale condition is naturally observed.
Apart from the failure of martingale property in Islah’s approximation, a fast and exact algorithm for sampling $F_{t+h}$ according to Eq. (20) has not been available in the literature. The methods presented so far are either slow (but accurate) or inaccurate (but fast). Chen et al. (2012) (partially) and Cai et al. (2017) adopted numerical root finding of $P_{\chi^{2}}(\,\cdot\,,\,\cdot\,,z_{K})=U$ for a uniform random number $U$, which belongs to the former type. This approach is slow due to the cumbersome evaluations of $P_{\chi^{2}}$ despite using some good initial guess and efficient calculation algorithm. Grzelak et al. (2019) speeded up sampling by using the stochastic collocation method, which is an optimal caching of the distribution function. Chen et al. (2012) (partially) and Cui et al. (2021) adopted the moment-matched quadratic Gaussian approximation to Eq. (20), which is proven to be effective in the Heston simulation (Andersen, 2008). While fast, it becomes inaccurate for a larger time step as it only uses the first two moments. This approach belongs to the latter type.
It is interesting to note that the exact CEV sampling algorithm in Section 4.2 can also be modified to sample $F_{t+h}$ under Islah’s approximation though it is not recommended due to the inherent drawbacks of Islah’s approximation. We outline the procedure in B. With the use of the exact CEV sampling procedure, we numerically illustrate the supermartingale property of $F_{t}$ in Section 5.
@@ 5 Numerical Results
We demonstrate the accuracy and effectiveness of our simulation algorithm with an extensive set of numerical examples. Table 1 shows the five parameter sets used in our numerical tests. For easier comparison to other existing methods, we select the sets from previous literature: the first two from Antonov and Spector (2012) and the rest from Cai et al. (2017). Following the literature (Chen et al., 2012; Cai et al., 2017), we price European options. Although the versatility of our simulation scheme goes beyond prcing European options, it offers a good testing case. Pricing European options also benefits from the availability of highly accurate benchmark option prices obtained from using the finite difference method (FDM) (Cai et al., 2017).

### v2

@@ 4.1 CEV approximation of the conditional forward price
$\bar{F}_{t}^{h}\approx F_{t}\exp\left(\frac{\rho(\sigma_{t+h}-\sigma_{t})}{\nu F_{t}^{\beta_{\ast}}}-\frac{\rho^{2}\sigma_{t}^{2}hI_{t}^{h}}{2F_{t}^{2\beta_{\ast}}}\right).$ (16b)
Algorithm 2 CEV approximation of the conditional terminal forward price
Note that our choice of $\bar{F}_{t}^{h}$ ensures the martingale property of conditional terminal forward price $F_{t+h}$. Based on Remark 1, we observe
@@
when $F_{t+h}$ is sampled according to Algorithm 2. The satisfaction of the martingale condition significantly improves the accuracy of our simulation algorithm (see numerical results illustrated in Figure 3).
4.2 Sampling conditional terminal forward price from a CEV distribution
With availability of analytic distribution function, the slow root-finding step is often used to invert the underlying distribution function. This approach has been adopted by some existing SABR simulation methods (e.g., Cai et al. (2017)). Our SABR simulation algorithm would be completed when our CEV approximation of the forward price in Algorithm 2 is paired with an efficient sampling algorithm of the CEV distribution. It is fortunate that we can use the algorithms of Makarov and Glew (2010) and Kang (2014) for the exact simulation of the CEV distribution. Despite its simplicity and efficiency, this exact simulation scheme of the CEV distribution is not widely known in the literature. First, we present the relevant probability distributions that are related to the CEV distribution and several related analytic properties.
@@ 4.2 Sampling conditional terminal forward price from a CEV distribution
Let ${\chi^{2}}(\delta,r)$ be the NCX2 random variable with degree of freedom $\delta$ and noncentrality $r$. The PDF and CDF of ${\chi^{2}}(\delta,r)$ are respectively given by
@@
The central chi-squared (corresponding to $r=0$) distribution is a gamma distribution. Let $\mathcal{G}(\alpha)$ be a gamma random variable with shape parameter $\alpha$ and unit scale parameter, the two random variables are related by
@@
The PDF and CDF of $\mathcal{G}(\alpha)$ are respectively given by  [^2] The CDF of $\mathcal{G}(\alpha)$, $P_{\mathcal{G}}(x;\alpha)$, is the scaled lower incomplete gamma function $\gamma(\alpha,x)$: $P_{\mathcal{G}}(x;\alpha)=\frac{\gamma(\alpha,x)}{\Gamma(\alpha)}\quad\text{where}\quad\gamma(\alpha,x)=\int_{0}^{x}t^{\alpha-1}e^{-t}\,\mathrm{d}t.$
@@
Definition 4 (Shifted Poisson variate).
The shifted Poisson (SP) variate with intensity $\lambda$ and shift parameter $\alpha$, $N\sim\mathcal{SP}(\lambda,\alpha)$, takes non-negative integer values with mass probability function:
@@
Next, we discuss the connection of the above defined distributions to the CEV distribution. It is well known that the transition density of $F_{T}$ is closely related to the NCX2 distribution (Schroder, 1989; Cox, 1996). We define $z_{T}$ be a transformation of the CEV process $F_{T}$ (see Eq. (8b)) by the function $z(\cdot)$:
$z_{T}:=z(F_{T})\quad\text{and}\quad z(y):=\frac{y^{2\beta_{\ast}}}{\beta_{\ast}^{2}\sigma_{0}^{2}T}.$ (21)
@@
Since the location variable $z$ plays the role of the noncentrality parameter, the distribution of $z_{T}$ is not a NCX2 distribution. The complementary CDF of $F_{T}>0$ and the probability of absorption (probability mass at zero) are respectively given by
$\mathrm{Prob}(F_{T}>y)=\mathrm{Prob}(z_{T}>z(y))=P_{\chi^{2}}\left(z_{0};\frac{1}{\beta_{\ast}},z(y)\right),$ (22a)
$\mathrm{Prob}(F_{T}=0)=\mathrm{Prob}(z_{T}=0)=1-P_{\chi^{2}}\left(z_{0};\,\frac{1}{\beta_{\ast}},0\right)=1-P_{\mathcal{G}}\left(\frac{z_{0}}{2};\,\frac{1}{2\beta_{\ast}}\right).$ (22b)
@@
Proposition 3 (Mixture gamma representation of the CEV distribution).
Under the CEV model with absorbing boundary condition, given $z_{T}>0$ ($F_{t}$ is not absorbed at the origin until time $T$), the transition from $z_{0}$ to $z_{T}$ can be sampled by an SP-mixture gamma distribution (Makarov and Glew, 2010, § 3.3):
$z_{T}\sim 2\,\mathcal{G}(N+1)\quad\text{where}\quad N\sim\mathcal{SP}\left(\frac{z_{0}}{2},\frac{1}{2\beta_{\ast}}\right).$ (23)
@@
$=P_{\mathcal{G}}\left(\frac{z_{0}}{2};\,\alpha\right)\sum_{k=0}^{\infty}P_{\mathcal{SP}}\left(k;\frac{z_{0}}{2},\alpha\right)\cdot\frac{1}{2}f_{\mathcal{G}}\left(\frac{z_{T}}{2};k+1\right),$
@@
As proposed by Kang (2014), the SP variable, $N\sim\mathcal{SP}(\lambda,\alpha)$, can be sampled by a gamma-mixture Poisson variable:
$N\sim\mathcal{POIS}(\lambda-X)\quad\text{where}\quad X\sim\mathcal{G}(\alpha)\quad\text{conditional on}\quad X<\lambda.$ (24)
Combining Eqs. (23) and (24) leads to an exact simulation algorithm for sampling the CEV distribution. Moreover, Scheme (24) can be further optimized in the context of the CEV simulation. This is the reason we select ‘Algorithm 3’ among the three algorithms in Kang (2014). As this advantage is not mentioned explicitly in Kang (2014), we state this feature as a remark:
@@
Suppose simulating $N\sim\mathcal{SP}(\lambda,\alpha)$ is the final goal, $X<\lambda$ can be obtained by repeatedly sampling $X$ until $X<\lambda$. If $N$ is an intermediate variable for the CEV simulation in Scheme (23), however, the repeated sampling of $X$ is unnecessary since
@@
Algorithm 3 Exact simulation of the CEV distribution of $F_{T}$ as depicted in Eq. (8b)
The CEV sampling algorithm can be applied to the conditional distribution of $F_{t+h}$ in Eqs. (16a) and (16b). We need to sample $z_{t+h}=z(F_{t+h})$ from $z_{t}=z(\bar{F}_{t}^{h})$, where the transformation $z(\cdot)$ is modified to become
@@
 $I_{t}^{h}\leftarrow$ Algorithm 1 ;
 $\bar{F}_{t}^{h}\leftarrow$ Eq. (16b) ;
@@
4.3 Comparison to Islah’s approximation used in other simulation schemes
While the earlier SABR simulation methods use various approaches for sampling $I_{t}^{h}$ (Simulation step 1), almost all of them (Chen et al., 2012; Cai et al., 2017; Leitao et al., 2017a, b; Grzelak et al., 2019; Cui et al., 2021; Kyriakou et al., 2023) are based on Islah’s approximation for sampling $F_{t+h}$ (Simulation step 2). We compare our approach to Islah (2009)’s in this section. We state Islah’s approximation as follows. For $0<\beta<1$, conditional on $\sigma_{t+h}$ and $I_{t}^{h}$ (and the filtration up to time $t$), the distribution of $F_{t+h}$ is approximated by
@@ 4.3 Comparison to Islah’s approximation used in other simulation schemes
$z^{\prime}_{t}=\frac{1}{\beta_{\ast}^{2}\rho_{\ast}^{2}\sigma_{t}^{2}hI_{t}^{h}}\left[F_{t}^{\beta_{\ast}}+\frac{\beta_{\ast}\rho}{\nu}(\sigma_{t+h}-\sigma_{t})\right]^{2}\quad\text{and}\quad z(y)=\frac{y^{2\beta_{\ast}}}{\beta_{\ast}^{2}\rho_{\ast}^{2}\sigma_{t}^{2}hI_{t}^{h}}.$
For the rationales behind this approximation, see Islah (2009, § 2.2), Chen et al. (2012, Result 2.4), or Cui et al. (2021, Appendix B.1).
Islah’s approximation is compared with the probability under our CEV approximation in Eqs. (16a) and (16b):
@@
The key difference between these two approximations is that Islah’s approximation uses the modified degree of freedom, $(1-\beta_{\ast}\rho^{2})/(\beta_{\ast}\rho_{\ast}^{2})$, in Eq. (25), and it is not equal to $1/\beta_{\ast}$ (except $\rho=0$). However, the transformation $z(\cdot)$ uses $\beta_{\ast}$ inconsistently. On the other hand, our CEV approximation uses $1/\beta_{\ast}$, which is the same degree of freedom for the CEV distribution. This means that $F_{t+h}$ under Islah’s approximation is not a CEV distribution, which leads to two drawbacks discussed in the next two remarks.
@@
In order to preserve the martingale property, Leitao et al. (2017a, b) artificially add an additive adjustment, commonly called the martingale correction, to the simulated samples of $F_{t+h}$. This ad-hoc martingale correction is inferior when compared with our approach where the martingale condition is naturally observed.
Apart from the failure of the martingale property in Islah’s approximation, a fast and exact algorithm for sampling $F_{t+h}$ according to Eq. (25) has not been available in the literature. The methods presented so far are either slow (but accurate) or inaccurate (but fast). Chen et al. (2012) (partially) and Cai et al. (2017) adopted numerical root finding of $P_{\chi^{2}}(\,\cdot\,,\,\cdot\,,z_{K})=U$ for a uniform random number $U$, which belongs to the former type. This approach is slow due to the cumbersome evaluations of $P_{\chi^{2}}$ despite using some good initial guess and efficient calculation algorithm. Grzelak et al. (2019) speeded up sampling by using the stochastic collocation method, which is an optimal caching of the distribution function. Chen et al. (2012) (partially) and Cui et al. (2021) adopted the moment-matched quadratic Gaussian approximation to Eq. (25), which is proven to be effective in the Heston simulation (Andersen, 2008). While being fast, it becomes inaccurate for a larger time step as it only uses the first two moments. This approach belongs to the latter type.
It is interesting to note that the exact CEV sampling algorithm in Section 4.2 can also be modified to sample $F_{t+h}$ under Islah’s approximation, though it is not recommended due to the inherent drawbacks of Islah’s approximation. We outline the simulation algorithm in B. With the use of the exact CEV sampling procedure, we numerically illustrate the supermartingale property of $F_{t}$ in Section 5.
@@ 5 Numerical Results
We demonstrate numerical accuracy and effectiveness of our simulation algorithm with an extensive set of numerical examples. In Subsection 5.1, we examine separately the errors caused by the two steps of simulation approximation: (i) SLN approximation of the average variance, (ii) CEV approximation of the forward price distribution. We then compare accuracy and speed tradeoff with analytic approximation formula pf European option prices in Subsection 5.2 and other simulation schemes in Subsection 5.3. In the last subsection, we show the competitive edge of accuracy performance of our CEV approximation of the forward price distribution with that of the Islah’s approximation.
@@
For easier comparison to other existing methods, we select the sets from the previous literature: the first two from Antonov and Spector (2012) and the rest from Cai et al. (2017). Following the literature (Chen et al., 2012; Cai et al., 2017), we price European options. Although the versatility of our simulation scheme goes beyond prcing European options, it offers a good testing case. Pricing European options also benefits from the availability of highly accurate benchmark option prices obtained from using the finite difference method (FDM) (Cai et al., 2017).
In our numerical tests, unless otherwise specified, we used $N=10^{5}$ paths for each simulation, and repeated $m=50$ times. Out of $m$ prices, we computed the average bias (used the FDM price as the benchmark) and standard deviation. Our algorithm is implemented in MATLAB R2023b on a laptop with a 13th Gen Intel Core i7-1360P 2.20 GHz processor. In Section 5.3, to demonstrate the speed of algorithm, we compare the CPU time of our method with those reported in Cai et al. (2017). Cai et al. (2017) used Matlab 7 on an Intel Core2 Q9400 2.66GHZ processor. We obtained the computer codes of Cai et al. (2017) and reran the codes again using our computer. The CPU times of their algorithm have been updated in Table 6 and Figure 2, where we examine the errors versus CPU time tradeoff. As shown in the comparison, our algorithm is seen to remain faster than the existing algorithms by at least one order of magnitude.
5.1 Examination of numerical errors caused by separate simulation steps
Recall that there are two separate simulation steps that may contribute to the overall errors in the full simulation scheme. The first step involves the SLN approximation of the average variance. The second step involves the CEV approximation of the forward price. The second step gives rise to errors from the operator splitting procedure and frozen coefficient approximation. The simulation of the CEV distribution via the SP mixture Gamma variate is exact, so this does not lead to simulation errors. We manage to isolate the errors that are separated into contributions from the SLN approximation and frozen coefficient approximation by considering accuracy performance of several degenerate cases of the SABR model. The accuracy analysis of the overall errors in the full simulation scheme in the general SABR model is discussed in Subsections 5.2 and 5.3.
First, we consider the lognormal SABR model, where $\beta=1$. The corresponding conditional forward price $F_{T}$ follows the lognormal distribution (see Eq. (6)), so one does not need to introduce operator splitting and frozen coefficient procedures to obtain an approximation of conditional $F_{T}$. The lognormal SABR model serves to isolate the error contributed by the SLN approximation only. We performed accuracy analysis of European option prices computed using our simulation scheme with the SLN approximation and the benchmark numerical results obtained from the finite difference method (FDM). We compute the relative error using the formula:
$\text{relative error}=\frac{\text{simulation calculation}-\text{FDM}}{\text{FDM}}.$
In our calculations, we use the following parameter values: $\beta=1$, $T=1$, $\nu=0.2$, $\rho=-0.75$ and $h=1$ for the base case, and perturb the parameter values to $\rho=-0.5$, $\rho=-0.25$, $\nu=0.4$, $\nu=0.8$, $h=0.5$ and $h=0.25$ as well. The other parameter values include $\sigma_{0}=0.2$ and $F_{0}=K=1$. The relative errors of the simulation calculations are listed in Table 1.
$\rho$ -0.75 -0.75 -0.75 -0.5 -0.25 -0.75 -0.75
$\nu$ 0.2 0.2 0.2 0.2 0.2 0.4 0.6
$h$ 1 0.5 0.25 1 1 1 1
Rel. Err. 0.00353% 0.00489% 0.0110% 0.00700% 0.00275% 0.00808% 0.0198%
Table 1: The relative errors contributed by the SLN approximation in the calculations of European call option prices for the lognormal SABR model ($\beta=1$). The FDM value for $\rho=-0.75$, $\nu=0.2$ is 0.07910, $\rho=-0.5$, $\nu=0.2$ is 0.07942, $\rho=-0.25$, $\nu=0.2$ is 0.07969, $\rho=-0.75$, $\nu=0.4$ is 0.07860 and $\rho=-0.75$, $\nu=0.6$ is 0.07811.
We observe that the errors contributed by the SLN approximation for the calculations of European option prices for the lognormal SABR model ($\beta=1$) are exceedingly small, only a very small fraction of one percentage point. In most cases, the agreement of the simulation result and FDM value is up to 3 significant figures. The errors increase with higher value of $\nu$, but they remain to be very small. The decrease of time step $h$ is irrelevant in decreasing the errors. Indeed, errors increase due to roundoffs with more calculation due to an increase of the number of time steps. In conclusion, the SLN approximation of the average variance is highly accurate and contributes almost insignificant errors.
Next, we consider the SABR models with $\rho=1$, $\rho=0.75$ and $\rho=0$ while $0<\beta<1$. The full correlation case ($\rho=1$) exhibits errors in calculating the European option prices contributed by the SLN approximation (should be exceedingly small from the above discussion) and frozen coefficient approximation, but no errors contributed from the operator splitting procedures. The conditional forward price of the zero correlation ($\rho=0$) SABR model follows the CEV distribution exactly, so there are no errors contributed by the operator splitting and frozen coefficient procedures. The errors arise only from the SLN approximation and simulation of the CEV distribution via the SP mixture Gamma variate. For $\rho=0.75$, it should have all three error sources: SLN approximation, operator splitting and frozen coefficient approximation. Table 2 lists the relative errors of the simulation calculations, which are benchmarked with the FDM results. Parameter values for $\sigma_{0}$, $F_{0}$, $K$ and $T$ are the same as those used in Table 1.
$\rho=1$ $\nu=0.2$ $\nu=0.2$ $\nu=0.2$ $\nu=0.4$ $\nu=0.8$
$\beta=0.4$ $\beta=0.6$ $\beta=0.8$ $\beta=0.8$ $\beta=0.8$
$h=1$ 0.518% 0.348% 0.164% 0.404 0.746%
$h=0.5$ 0.418% 0.225% 0.0683% 0.199 0.350%
$h=0.25$ 0.244% 0.119% 0.0299% 0.0947 0.224%
$\rho=0.75$ $\nu=0.2$ $\nu=0.2$ $\nu=0.2$ $\nu=0.4$ $\nu=0.8$
$\beta=0.4$ $\beta=0.6$ $\beta=0.8$ $\beta=0.8$ $\beta=0.8$
$h=1$ 0.415% 0.306% 0.125% 0.333% 0.421%
$h=0.5$ 0.237% 0.130% 0.0647% 0.0773% 0.301%
$h=0.25$ 0.0399% 0.0287% 0.0242% 0.0597% 0.215%
$\rho=0$ $\nu=0.2$ $\nu=0.2$ $\nu=0.2$ $\nu=0.4$ $\nu=0.8$
$\beta=0.4$ $\beta=0.6$ $\beta=0.8$ $\beta=0.8$ $\beta=0.8$
$h=1$ -0.0562% -0.00574% 0.0704% 0.0257% 0.123%
$h=0.5$ -0.0138% 0.0151% -0.0537% 0.0803% 0.0292%
$h=0.25$ -0.0526% 0.0123% -0.0162% 0.0616% 0.0758%
Table 2: The relative errors in calculating the European call option prices for the full correlation ($\rho=1$), $\rho=0.75$ and zero correlation ($\rho=0$) SABR models. The FDM value for $\rho=1$: $\nu=0.2\,,\beta=0.4$ is 0.07989, $\nu=0.2,\,\beta=0.6$ is 0.08002, $\nu=0.2,\,\beta=0.8$ is 0.08017, $\nu=0.4,\,\beta=0.8$ is 0.08044 and $\nu=0.8,\,\beta=0.8$ is 0.08043. The FDM values for $\rho=0.75$ are 0.07998, 0.08008, 0.08018, 0.08083 and 0.08276, respectively. The FDM values for $\rho=0$ are 0.07996, 0.07994, 0.07992, 0.08068 and 0.08355, respectively.
For the full correlation ($\rho=1$) SABR model, the main source of errors comes from the frozen coefficient approximation. In all cases of different values of $\nu$ and $\beta$, the errors show decrease in magnitude with respect to the time step $h$. The frozen coefficient approximation is sensitive to the size of the time step $h$, smaller approximation error with smaller time step. Also, the errors show increase in magnitude with respect to $\beta_{\ast}=1-\beta$ and $\nu$. Such trends hold similarly for $\rho=0.75$, whose errors are from the three main sources: SLN approximation, operator splitting and frozen coefficient approximation. It is seen that the errors from operator splitting and frozen coefficient approximation may counteract with each other, leading to a smaller error when compared with the full correlation ($\rho=1$). On the other hand, for zero correlation ($\rho=0$) SABR model, the sources of errors come from the SLN approximation and exact simulation of the CEV distribution, both are exceedingly small (well below one percentage point). There is no pattern of dependence on $\nu$, $\beta$ and $h$ since errors are mostly from roundoff.
5.2 Comparison to analytic approximations